# Chen Feng1 Duolikun Danier1 Fan Zhang David Bull

### 14 Dec 2023 v1

@@ I Introduction
The rest of the paper is organised as follows. Section II describes the collection/creation of source sequences, the generation of each visual artefact and the annotation of artefact labels. Section III presents the benchmarked detection methods and how their performance is evaluated. The detection results are then shown in Section IV alongside the analysis and discussion. Finally, Section V concludes the paper and outlines future works.
@@ II-B Source Artefacts
Motion blur typically results from insufficient frame rates or relatively wide shutter angles, with areas corresponding to the rapidly moving parts of the scene with reduced high frequency energy. While motion blur can improve the perception of motion, it can also degrade visual quality when excessive [11]. Considering the difficulty of synthesising realistic motion blur, ten source videos are collected with this artefact.
Dark scenes, when described as artefacts, reflect insufficient lighting (due to underexposure during acquisition) that leads to decreased perceptual quality. Although physics-based lighting degradation models [18, 19] exist for synthesising dark content, they generally yield reduced realism; we therefore selected 10 natural dark sequences for this database.
@@ II-D Artefact Label Annotation
For each of the 480 sequences, we provide a binary label for each of the ten artefact types, indicating their presence. These labels were first annotated according to the artefact generation methods described above, and then confirmed through manual inspection to ensure that the labelled artefacts are indeed recognisable. As a result, for each source artefact, there are 80 (i.e. 16.7%) positive videos (i.e. the artefact is visible), and for each non-source artefact, there are around 240 (i.e. 50%) positive samples. Additionally, for some non-source artefacts (i.e. spatial blur, dropped frames, black frames), we also provided frame-level binary annotation.

### 07 Mar 2024 v2

@@ II-B Source Artefacts
Motion blur typically results from insufficient frame rates or relatively wide shutter angles, with areas corresponding to the rapidly moving parts of the scene with reduced high frequency energy. While motion blur can improve the perception of motion, it can also degrade visual quality when excessive [11]. Considering the difficulty of synthesising realistic motion blur, ten source videos are collected with this artefact with various blur levels.
Dark scenes, when described as artefacts, reflect insufficient lighting (due to underexposure during acquisition) that leads to decreased perceptual quality. Although physics-based lighting degradation models [18, 19] exist for synthesising dark content, they generally yield reduced realism; we therefore selected 10 natural dark sequences for this database with different darkness levels.
@@ II-D Artefact Label Annotation
For each of the 480 sequences, we provide a binary label for each of the ten artefact types, indicating their presence. These labels were first annotated according to the artefact generation methods described above, and then confirmed through manual inspection to ensure that the labelled artefacts are indeed recognisable. As a result, for each source artefact, there are 80 (i.e. 16.7%) positive videos (i.e. the artefact is visible) [^1] This does not affect the use of this database for source artefact detection, as described in Section II-E, and for each non-source artefact, there are around 240 (i.e. 50%) positive samples. Additionally, for some non-source artefacts (i.e. spatial blur, dropped frames, black frames), we also provided frame-level binary annotation.